OpenStudy (anonymous):

please help √3 (tan255-tan195)

OpenStudy (aravindg):

Try to express these angles as sum For eg: For tan 225 tan(255) = tan(180+75) = tan 75 = tan(30 + 45) Then use ta(A+B) formula